Page tree
Skip to end of metadata
Go to start of metadata

10 Attendees

Apologies

  • Martin Kolleck (PHI)

Agenda

Discussion items

ItemNotes
Contact re VM deliveries etc.

RC explained that to avoid any delays in the case that he is not available (e.g. on leave), it is best for the instrument teams to email solarorbiter_sciops@sciops.esa.int to notify the SOC, for example, that they have uploaded a new VM version for testing.
AW clarified that this is the same address being used to consolidate other communication with the SOC relating to operations.

VM Status - engineering.

FE asked that RC check if log rotation now seen to be working.

GL mentioned that Chandra will soon (this week or next) be looking to the problem causing segmentation faults inside the SWA VM (including 00.07.00), and at making the VM software more robust so as to carry on after such seg faults.

LH mentioned that the only pending Solo-HI update was to modify certain header values, for which she was waiting for input from scientists.

Maurizio has not recently had time to work on the VM. In commissioning a bug was found in the on-board application software that it does not pass on the start and stop times for the light-curves. They are currently attempting to find a workaround using the time of the first packet. This will have to be incorporated in the VM.

For SPICE Terje mentioned that not much change in the last few months. He has fixed a few bugs in the science pipeline and should update the VM to incorporate these fixes. He may do this by the end of August. But there is nothing critical.

For MAG Vincent said he had no updates ready yet, but he was aware that log rotation and should-fail tests were pending. He hopes to have something on these by the next telecon.

During the meeting we verified that the current operationally deployed STIX VM is 00.06.09. And the SOC has a version 00.07.01. RC acknowledged that he still had to restart the test VM after deleting the old log files to see if log rotation then worked.

VM product status

AW asked if the SWA v7 VM includes HIS Helium moments. GL answered (later during AoB) that yes it does.

Terje wanted to know if the SOC could clarify an allusion by Dave to an ambiguity in a descriptor. Anik said that she would look for any reference to it in SOC internal documents.

AoB

VE wanted to know when the LL data will go live in the archive, and whether MAG will get a chance to review the contents beforehand.
Andrew agreed to create a zip and send it for review.

VE also mentioned that a new calibration bundle would be delivered soon.

GL wanted to know which web address to use to access the archive. AW said that the SOAR address would be circulated at the time of release. The Low Latency Display Tool will come later.

We decided that the next telecon should be on the 28th of October.

Preparatory Actions

New Actions

None.